OpenStudy (anonymous):

What is the solution to 3^4x-7= 4^2x+3

OpenStudy (anonymous):

Is it \[3^{4}x - 7 = 4 ^{2}x + 3\] or \[3^{4x} - 7 = 4 ^{2x} + 3\]

sam (.sam.):

or \[3^{4x-7}=4^{2x+3}\]

OpenStudy (anonymous):

Sam's right

OpenStudy (anonymous):

10/65

OpenStudy (anonymous):

explanations please?

OpenStudy (anonymous):

\[3^{4x-7}=4^{2x+3}\] \[\frac{3^{4x}}{3^{7}}=4^{2x}\times {4^3}\] \[\frac{81^x}{2187} = 64\times16^x\] \[(\frac{81}{16})^x = 64 \times 2187\] Now takes logs on both sides

OpenStudy (anonymous):

\[x = \frac{(\log 64+\log2187)}{\log 81-\log 16}\]

OpenStudy (anonymous):

wow thanks!